OpenStudy (anonymous):

the solution to log(x+9)=log4x is equal to 4. true or false?

OpenStudy (anonymous):

graph

OpenStudy (owlcoffee):

Let's check it: \[\log_{}(x+9)=\log_{} 4x\] Now, this is exactly the same as the exponential equations. If we hace the same logaritms we can make everything a polinomial equality with the arguement of the logarithm. so by that statement I can say this: \[x-9=4x\] sustracting 4x to both sides, and adding 9 as well we end up with: \[-3x=9\] \[x=\frac{ 9 }{ -3 }\] \[x=-3\] So with that,we can say that the statement is false.

OpenStudy (ranga):

Is x = 4 the solution to log(x+9) = log4x ? Put x = 4. LHS is log(13), RHS is log(16). They are not the same and so x = 4 is not a solution.

OpenStudy (anonymous):

thank yall!